OpenStudy (anonymous):

Find dy/dx: x^(2)y-2y+5=0 Find dy/dx: x^(2)y-2y+5=0 @Mathematics

OpenStudy (turingtest):

\[{d \over dx}(x^2y-2y+5)=0\]\[2xy+x^2y'-2y'=2xy+y'(x^2-2)=0\]\[y'={2xy \over 2-x^2}\]

OpenStudy (anonymous):

Oh i got a simular answer